OpenStudy (anonymous):

The standard acceleration of gravity at the earth’s surface is 9.81m/s^2. Convert this acceleration to units of ft/h^2 and mi/s^2

OpenStudy (anonymous):

@bibby Please help me! I even posted the question in the box ._.

OpenStudy (bibby):

you're making progress *~* <3

OpenStudy (bibby):

for units squared write them out as s*s like so

OpenStudy (anonymous):

I know that for the mi/s^2 the answer is 0.0060956513958482 mi/s2 but I don't know how though

OpenStudy (bibby):

ft/h^2\(\huge \frac{9.81m}{1\cancel{s}*1\cancel{s}}*\frac{60\cancel{s}*60\cancel{s}}{1\cancel{min}*1\cancel{min}}*\frac{60\cancel{min}*60\cancel{min}}{1hr*1hr}\)

OpenStudy (anonymous):

I got a really big number though I got 127137600 ft/h^2

OpenStudy (bibby):

well that's meters per hour^2, now we have to get it into feet

OpenStudy (bibby):

you're not noticing the units being cancelled

OpenStudy (anonymous):

we just need ft now because we already have the hr^2 right?

OpenStudy (bibby):

yeah, we have m/hr^2

OpenStudy (anonymous):

so now would it be 127137600m/hr^2 divided by 1000m?

OpenStudy (bibby):

idk how many meters are in a foot?

OpenStudy (anonymous):

hold up let me search it

OpenStudy (anonymous):

there is .3048 meters in a foot

OpenStudy (anonymous):

@dumbcow @nincompoop @inkyvoyd help please?

OpenStudy (anonymous):

I NEED HELP I REALY DO SOMEBODY PLEASE HELP ME IM SERIOUSLY ABOUT TO SHOOT MYSELF I JUST CANT ANYMORE. I BEEN HERE SINCE 11 AND SO FAR I HAVENT GOTTEN ANY REAL HELP JUST STUPID COMMENTS FROM MR.owlet -_-